Shift Summary for owl shift, 01/12/2000


Subject: Shift Summary for owl shift, 01/12/2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:20:53 CST


              Shift Summary for owl shift, 01/12/2000

On Shift: J. Shields, A. Tripathi

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 7.0

Accounting of lost time:

========================================================================
Runs ended this shift:

Run 15484 -- SHORT_REG832

Tapes brought to FCC this shift:
  (including tapes for online split)

QKE058-073, QKB154-155, UPH057, UPX020, UPLA15

========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:

Special Runs: Short Regenerator, SEM~2.4e12. Numerous KTEVANA/pipeline
  errors.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15484 00:00 inherit on-going run from previous shift

      ~237 00:37 ring counters removed from readout for a spill
                       or two.

              01:30 R.K makes some adjustment to improve livetime:
                        set REG ADC thresholds to 255 for un_used channels,
                        set UCAL test ADC (11 bit) threshold to 255 counts
                        instead of zero, set RC ADC thresholds to 255
                        instead of zero.

              02:06 R.K. adds kfrend_832_noadc_txt script ( fast kfrend w/o
                       Fera ADC resets) to the kfrend options

       313 02:17 BA ADC gate delayed 16 ns for UCAL test. Valid
                       for all spills > 313

              02:30 L3 accept fraction error. Run stopped,
                       did drt_fresh_start, and changed tapes.
                       Mounted QKE074-082, QKB156

15485 1 02:45 started run 15485. Expected changes from previous
                       run: x10 B05 triggers for monitoring, 50% less
                       B02 triggers for CsI (PS = 15 instead of 9), and
                       minor changes to HDIB trigger
                     
              03:45 R.K turns off HDIB triggers for a few spills
                       and the 2V source timing looks better --> triggers
                       w/o REG cause strange 2V shape. Note that 70%
                       of triggers to PL3 don't use REG

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:20:55 CST